Identify the error in the problem.
Solve 5y-10x=15 for y.
Step 1: 5y=15+10x
Step 2: y=15+10x/5
Step 3: y=3+10x

Found 2 solutions by josgarithmetic, MathTherapy:
Answer by josgarithmetic(39616)

Solve 5y-10x=15 for y.

ADDITION PROP of EQLTY, Step 1: 5y=15+10x

MULTIPLICATION PROP of EQLTY, Step 2: y=15+10x/5
Mistake in grouping

DISTRIBUTIVE PROPERTY, Step 3: y=3+10x
The needed grouping WAS observed, but one of the computational parts was done wrongly.

The correct result should be y=3+2x.

Answer by MathTherapy(10551)


Identify the error in the problem.
Solve 5y-10x=15 for y.
Step 1: 5y=15+10x
Step 2: y=15+10x/5
Step 3: y=3+10x
Step 2 error: 15 on the right also needs to be divided by 5, not just 10x on that side